# A Survey on Data Selection for Language Models

### 26 Feb 2024 v1

@@ 8 Data Selection for Other Domains
Of particular interest has been the rise in pretrained vision-language models, which has followed closely after the success of large language models (Radford et al., 2021; Schuhmann et al., 2022; Zhu et al., 2023b). For example, Schuhmann et al. (2022) and Zhu et al. (2023b) design datasets for vision-language pretraining using some of the same methods and ideas presented in §3 as well as some methods that are specific to their vision-language data. Schuhmann et al. (2022) collect web-text along with images that contain an alt-text in order to create data with image-text pairs. They utilize some heuristic filters on the text, similar to those from §3, but they also remove data with less than 5kb of image data, as well as image-text pairs that have cosine similarity less than 0.28 for english data and 0.26 for other languages (image and text separately encoded using OpenAI’s ViT-B/32 CLIP model). The similarity step removed significant quantities of data, reducing the original corpus from around 50 billion images to 6 billion. While LAION-5B is a corpus entirely of image-text pairs, Multimodal C4 (Zhu et al., 2023b) introduces a dataset of interleaved sentences and images, where each data point can contain multiple sentences and images. Zhu et al. (2023b) use a preprocessed English-only C4 dataset and gather image files (png/jpeg/jpg) that were found on the source websites for C4. They perform an image-deduplication step to ensure that a single image is not overrepresented within a single document or across the entire corpus. Then, they discard images with a height or width smaller than 150 pixels, which accounts for many small icons and navigation buttons, as well as images with aspect ratios greater than 2 or less than 0.5 (removes many banners at the top/bottom or sides of websites). Finally, to interleave images into the existing text, they use cosine similarity between images and individual texts, interleaving only the images which have a similarity score of at least 0.15. Another unique filtering done for this domain is a step of removing images that contain faces.
@@
These methods mainly filter data with heuristic approaches, and have the same drawbacks as heuristic approaches for language datasets; namely, the difficulty of validating a method, and slow iteration time as it requires training multiple models. Data Filtering Networks aim to overcome these weaknesses by filtering data not with heurstics, but with a trained network (Fang et al., 2023). While Fang et al. (2023) demonstrate that it is possible to train a filtering network that can select data points which lead to a better model, in a cruel twist, they find that data quality is key to training a good filtering network, leading to a circular problem.
In addition to data collection and filtering, data mixing is also a topic of concern for non-language pretraining. Piergiovanni et al. (2023) propose an online data mixing method for vision-language models that mixes a variety of tasks together (e.g. image captioning or visual question answering). Their method, dynamic difficulty sampling, calculates the loss of each task $t$, $\mathcal{L}_{t}$, and weights each task proportionally with the total loss $L$, getting the final sampling weight as $S_{t}=\dfrac{\mathcal{L}_{t}}{L}$. This mixing method maintains some similarities to the online data mixing methods from language model pretraining (Xia et al., 2023a; Albalak et al., 2023a; Schioppa et al., 2023).
@@
A number of methods have been developed to improve the data selection on the level of individual batches, sometimes called data difficulty estimation. One of the motivations of these methods is that they use the current model’s state to determine which data to select, but this can add significant computational overhead. Additionally, because these methods are run alongside the model training, their cost is not amortized across training runs as it would be for static datasets. Chang et al. (2017) propose two methods for Active Bias that emphasize training on data points where the model is uncertain, based on their prediction variance or closeness to the decision threshold. Similarly, Song et al. (2020) propose Recency Bias, a method that further emphasizes training on uncertain data points that were predicted inconsistently in recent iterations of training. Finally, (Mindermann et al., 2022) present Reducible Holdout Loss Selection (RHO-Loss), a method of adaptive batch selection that, intuitively, avoids selecting data points that are redundant, noisy, or less relevant. RHO-Loss was validated on a very small sample of NLP tasks, task-specific fine-tuning on SST2 (Socher et al., 2013) and CoLA (Warstadt et al., 2019), but has not been demonstrated to work in other learning settings.
@@ 9 Related Topics
In contrast, the objective of coreset selection (Mirzasoleiman et al., 2020a; Borsos et al., 2020; Killamsetty et al., 2021b; Guo et al., 2022; Killamsetty et al., 2021a) is to select a subset from the given dataset so that a model trained on it achieves optimal performance, which better aligns with the topics discussed in this paper. Many coreset selection methods design a series of scores and rank the data instances accordingly. Within this realm, geometry-based coreset selection methods selects the examples based on their distances from k-means cluster centers (Chen et al., 2012; Sener & Savarese, 2018; Sorscher et al., 2022; Xia et al., 2023b). Additionally, predictive uncertainty and model loss metrics can also serve as criteria for ranking data instances (Coleman et al., 2019; Pleiss et al., 2020; Paul et al., 2021; Zheng et al., 2022).
There also exist coreset selection methods that are optimization-driven. Craig (Mirzasoleiman et al., 2020a) optimizes the coreset selection so that the sum of the gradients on the subsets closely match the gradients on the full dataset. This approach has been extended to handle noisy data settings in the subsequent work (Mirzasoleiman et al., 2020b). Glister (Killamsetty et al., 2021b) and Grad-Match (Killamsetty et al., 2021a) focus on coreset selection on noisy and imbalanced datasets and propose to match the gradient of the coreset with the training of a held-out validation set. The AdaCore (Pooladzandi et al., 2022) method incorporates the second-order gradient information for better gradient matching. Notably, the techniques of coreset selection have found application in LLM fine-tuning (Liu et al., 2023b; Xia et al., 2024). For example, the LESS (Xia et al., 2024) method selects influential data for supervised fine-tuning based on the similarity of gradients evaluated on the coreset and the validation set.
@@
Data augmentation methods are used to generate new data by modifying the existing data points in a dataset. Similar to data selection, data augmentation also aims to shift the training distribution. Data selection is often used to improve the data coverage in domains where there is limited data (Li et al., 2022b; Chen et al., 2023a) and can be used in conjunction with existing data selection methods (e.g. domain-specific selection in §3.4 or “selecting from auxiliary data to improve generalization” in §7). For example, Albalak et al. (2022a) use a policy gradient along with information from an auxiliary task to improve their main model. However, the quality of the coverage produced from augmentation may be limited, as the semantics of augmented data can diverge from the underlying distribution of language, leading to an undesirable distribution shift and possibly a decrease in performance (Li et al., 2022b; Chen et al., 2023a). Thus, care should be taken when performing data augmentation to ensure that the augmentations do not stray too far from the desired distribution. Nonetheless, data augmentation can still be a useful tool, and we refer readers to existing surveys on the many methods for augmentation in NLP (Chen et al., 2023a; Feng et al., 2021) and computer vision (Yang et al., 2023).
@@
Curriculum learning has found limited success in the natural language domain. For example, Graves et al. (2017) use multi-armed bandits to determine an appropriate curriculum for language modelling with LSTMs. Xu et al. (2020) demonstrate that curricula can be used to improve learning in natural language understanding tasks for a BERT model. More recently, Fan & Jaggi (2023) propose irreducible curriculum for language model pretraining, which prioritizes training on samples with higher learnability (motivated by RHO-Loss (Mindermann et al., 2022)). Curriculum learning is one aspect of data-centric AI that has not yet seen a large adoption in large language models, in part due to its limited success.
@@ 10.1 Test set decontamination
Understanding a model’s capabilities requires evaluation on a fair benchmark, but the training of language models on web-scale data has raised concerns and speculation about the possible inclusion of evaluation benchmarks in training data. Given the scale of data it is not feasible to manually check, so automated methods have been designed to detect the existence of evaluation data in the training set (Rae et al., 2022; Li et al., 2022a; 2023a; Magnusson et al., 2023; Elazar et al., 2023), as well as detecting models that have already been trained on evaluation data (Carlini et al., 2018; Jagielski, 2023; Marone & Durme, 2023). Jacovi et al. (2023) on the other hand, suggest three strategies for reducing the risks of contamination. Data decontamination is an important step in ensuring the integrity of model evaluation, and is included in the discussion section because this critical step is relevant, and should be required, for all stages of model training.
@@
Test set contamination methods often follow a similar form to deduplication methods. When decontaminating MassiveTextRae et al. (2022) compute the 13-gram Jaccard similarity between train and test documents (as originally proposed by Lee et al. (2022a)) and remove train documents that have a Jaccard similarity over 0.8, the same exact method that they use to detect duplicated documents. Similarly, Brown et al. (2020) and Gao et al. (2020) use 13-gram overlap filtering to detect contamination in their respective training sets. To make decontamination easier, some tools have been developed. For example, CarperAI create a github repository with code for decontaminating evaluation data from a training dataset using MinHashLSH. [^20] CarperAI decontamination github repository Additionally, Marone & Durme (2023) propose Data Portraits, a method of investigating the data that a model was trained on. Data portraits are an artifact that records training data, allows downstream inspection, and enables answering questions about test set leakage and model plagiarism through a novel membership testing tool, a strided bloom filter. In addition to the new method, they release a demo of their tool [^21] https://dataportraits.org/ which uses 50-character grams to detect overlap between training data and an input sequence.
@@
Proving that black-box models have trained on a data point can be very challenging, but also important to ensure that their results can be trusted. Because their data is not disclosed, it is impossible to know with certainty what their training data was, but some methods have been proposed as efforts to address this issue. Shi et al. (2023a) propose min-kprob, using the intuition that unseen examples are more likely to contain outlier words with low probabilities according to the model, whereas unseen examples are less likely to exhibit such low probability words. While this is not explicitly a method of decontamination, it is a useful tool for determining how contaminated a pretrained models data was. This tool can be used for benchmark example contamination detection, privacy auditing for machine unlearning, and copyrighted text detection. Oren et al. (2023) present another method for detecting contaminated data based on the observation that models have a tendency to memorize the order of examples as they exist in the dataset, and provide provable guarantees of test set contamination through a framework of statistical tests.
@@
Decontamination in code data is particularly important, as Li et al. (2022a) suggest that competitors in coding competitions tend to publish their solutions online after a competition. One solution to this problem is to use a temporal split when dividing data into training and evaluation. To ensure that test data does not exist in the training dataset, the training dataset can be constructed only from files that originate prior to the test competition. This ensures that the training data includes only information that would be available for human participants of the competition. In a more traditional method, StarCoder (Li et al., 2023a) decontaminates The Stack by removing files that contained either exact matching docstrings or solutions from the HumanEval (Chen et al., 2021) or MBPP (Austin et al., 2021) datasets, docstrings from APPS (Hendrycks et al., 2021), questions from GSM8K (Cobbe et al., 2021), or prompts from DS1000 (Lai et al., 2023).
@@ 11.1 Accelerating research on data selection
Research on data can be a very slow and tedious process, potentially requiring significant investments of time, compute, and human effort. Here we describe three directions of future research that can reduce the effort required to do research in data selection.
@@
Preliminary works have been done in this direction, including some works on data attribution and valuation (Ghorbani & Zou, 2019; Ilyas et al., 2022), and data measurements (Mitchell et al., 2023). Works on data valuation and attribution (discussed in §9) attempt to better understand how each data point affects a models predictions, but may be used as inspiration for methods that directly evaluate the data itself. Additionally, there are a number of metrics that directly measure the intrinsic characteristics of data including distance, density, diversity, tendencies (mean, median, mode), and association (correlation, mutual information), but these measures have yet to be connected with a model’s performance on downstream data (Mitchell et al., 2023).
@@ 11.3 Shifting compute time from model training to data processing
One example of how this may be accomplished is through the use of methods which have previously only been used in low-data settings (e.g. alignment) and applying them to settings with greater amounts of data (e.g. pretraining). For example, methods inspired by model-based filtering, or reward model-reweighting techniques can be used to improve the “quality” of pretraining data. As compute gets cheaper, it becomes more feasible to spend the cost of directly applying the existing methods. However, another avenue worth exploring is using cheap approximations for utility functions, such as computing similarity using bag-of-n-grams and hash functions rather than from model representations.
@@ Acknowledgments
This material is based on work that is partially funded by an unrestricted gift from Google. This work was supported by the National Science Foundation award #2048122. The views expressed are those of the authors and do not reflect the official policy or position of the US government. The work of BH and SC is partially supported by the UC Santa Barbara IEE IGSB SW Impact Grant.

### 08 Mar 2024 v2

@@ 8 Data Selection for Other Domains
Of particular interest has been the rise in pretrained vision-language models, which has followed closely after the success of large language models (Radford et al., 2021; Schuhmann et al., 2022; Zhu et al., 2024). For example, Schuhmann et al. (2022) and Zhu et al. (2024) design datasets for vision-language pretraining using some of the same methods and ideas presented in §3 as well as some methods that are specific to their vision-language data. Schuhmann et al. (2022) collect web-text along with images that contain an alt-text in order to create data with image-text pairs. They utilize some heuristic filters on the text, similar to those from §3, but they also remove data with less than 5kb of image data, as well as image-text pairs that have cosine similarity less than 0.28 for english data and 0.26 for other languages (image and text separately encoded using OpenAI’s ViT-B/32 CLIP model). The similarity step removed significant quantities of data, reducing the original corpus from around 50 billion images to 6 billion. While LAION-5B is a corpus entirely of image-text pairs, Multimodal C4 (Zhu et al., 2024) introduces a dataset of interleaved sentences and images, where each data point can contain multiple sentences and images. Zhu et al. (2024) use a preprocessed English-only C4 dataset and gather image files (png/jpeg/jpg) that were found on the source websites for C4. They perform an image-deduplication step to ensure that a single image is not overrepresented within a single document or across the entire corpus. Then, they discard images with a height or width smaller than 150 pixels, which accounts for many small icons and navigation buttons, as well as images with aspect ratios greater than 2 or less than 0.5 (removes many banners at the top/bottom or sides of websites). Finally, to interleave images into the existing text, they use cosine similarity between images and individual texts, interleaving only the images which have a similarity score of at least 0.15. Another unique filtering done for this domain is a step of removing images that contain faces.
@@
These methods mainly filter data with heuristic approaches, and have the same drawbacks as heuristic approaches for language datasets; namely, the difficulty of validating a method, and slow iteration time as it requires training multiple models. Data Filtering Networks aim to overcome these weaknesses by filtering data not with heurstics, but with a trained network (Fang et al., 2024). While Fang et al. (2024) demonstrate that it is possible to train a filtering network that can select data points which lead to a better model, in a cruel twist, they find that data quality is key to training a good filtering network, leading to a circular problem.
In addition to data collection and filtering, data mixing is also a topic of concern for non-language pretraining. Piergiovanni et al. (2023) propose an online data mixing method for vision-language models that mixes a variety of tasks together (e.g. image captioning or visual question answering). Their method, dynamic difficulty sampling, calculates the loss of each task $t$, $\mathcal{L}_{t}$, and weights each task proportionally with the total loss $L$, getting the final sampling weight as $S_{t}=\dfrac{\mathcal{L}_{t}}{L}$. This mixing method maintains some similarities to the online data mixing methods from language model pretraining (Xia et al., 2024a; Albalak et al., 2023a; Schioppa et al., 2023).
@@
A number of methods have been developed to improve the data selection on the level of individual batches, sometimes called data difficulty estimation. One of the motivations of these methods is that they use the current model’s state to determine which data to select, but this can add significant computational overhead. Additionally, because these methods are run alongside the model training, their cost is not amortized across training runs as it would be for static datasets. Chang et al. (2017) propose two methods for Active Bias that emphasize training on data points where the model is uncertain, based on their prediction variance or closeness to the decision threshold. Similarly, Song et al. (2020) propose Recency Bias, a method that further emphasizes training on uncertain data points that were predicted inconsistently in recent iterations of training. Finally, (Mindermann et al., 2022) present Reducible Holdout Loss Selection (RHO-Loss), a method of adaptive batch selection that, intuitively, avoids selecting data points that are redundant, noisy, or less relevant. RHO-Loss was validated on a very small sample of NLP tasks, task-specific fine-tuning on SST2 (Socher et al., 2013) and CoLA (Warstadt et al., 2019), but has not been demonstrated to work in other learning settings, and Kaddour et al. (2024) show that it is not computationally efficient.
@@ 9 Related Topics
In contrast, the objective of coreset selection (Mirzasoleiman et al., 2020a; Borsos et al., 2020; Killamsetty et al., 2021b; Guo et al., 2022; Killamsetty et al., 2021a) is to select a subset from the given dataset so that a model trained on it achieves optimal performance, which better aligns with the topics discussed in this paper. Many coreset selection methods design a series of scores and rank the data instances accordingly. Within this realm, geometry-based coreset selection methods selects the examples based on their distances from k-means cluster centers (Chen et al., 2012; Sener & Savarese, 2018; Sorscher et al., 2022; Xia et al., 2023). Additionally, predictive uncertainty and model loss metrics can also serve as criteria for ranking data instances (Coleman et al., 2019; Pleiss et al., 2020; Paul et al., 2021; Zheng et al., 2022).
There also exist coreset selection methods that are optimization-driven. Craig (Mirzasoleiman et al., 2020a) optimizes the coreset selection so that the sum of the gradients on the subsets closely match the gradients on the full dataset. This approach has been extended to handle noisy data settings in the subsequent work (Mirzasoleiman et al., 2020b). Glister (Killamsetty et al., 2021b) and Grad-Match (Killamsetty et al., 2021a) focus on coreset selection on noisy and imbalanced datasets and propose to match the gradient of the coreset with the training of a held-out validation set. The AdaCore (Pooladzandi et al., 2022) method incorporates the second-order gradient information for better gradient matching. Notably, the techniques of coreset selection have found application in LLM fine-tuning (Liu et al., 2023b; Xia et al., 2024b). For example, the LESS (Xia et al., 2024b) method selects influential data for supervised fine-tuning based on the similarity of gradients evaluated on the coreset and the validation set.
@@
Data augmentation methods are used to generate new data by modifying the existing data points in a dataset. Similar to data selection, data augmentation also aims to shift the training distribution. Data selection is often used to improve the data coverage in domains where there is limited data (Li et al., 2022b; Chen et al., 2023a) and can be used in conjunction with existing data selection methods (e.g. domain-specific selection in §3.4 or “selecting from auxiliary data to improve generalization” in §7). For example, Albalak et al. (2022a) use a policy gradient along with information from an auxiliary task to improve their main model. However, the quality of the coverage produced from augmentation may be limited, as the semantics of augmented data can diverge from the underlying distribution of language, leading to an undesirable distribution shift and possibly a decrease in performance (Li et al., 2022b; Chen et al., 2023a). Thus, care should be taken when performing data augmentation to ensure that the augmentations do not stray too far from the desired distribution. Nonetheless, data augmentation can still be a useful tool, and we refer readers to existing surveys on the many methods for augmentation in NLP (Chen et al., 2023a; Feng et al., 2021), code (Zhuo et al., 2023), and computer vision (Yang et al., 2022).
@@
Curriculum learning has found limited success in the natural language domain. For example, Graves et al. (2017) use multi-armed bandits to determine an appropriate curriculum for language modeling with LSTMs. Xu et al. (2020) demonstrate that curricula can be used to improve learning in natural language understanding tasks for a BERT model. More recently, Fan & Jaggi (2023) propose irreducible curriculum for language model pretraining, which prioritizes training on samples with higher learnability (motivated by RHO-Loss (Mindermann et al., 2022)). Curriculum learning is one aspect of data-centric AI that has not yet seen a large adoption in large language models, in part due to its limited success.
@@ 10.1 Test set decontamination
Understanding a model’s capabilities requires evaluation on a fair benchmark, but the training of language models on web-scale data has raised concerns and speculation about the possible inclusion of evaluation benchmarks in training data. Given the scale of data it is not feasible to manually check, so automated methods have been designed to detect the existence of evaluation data in the training set (Rae et al., 2021; Li et al., 2022a; 2023b; Magnusson et al., 2023; Elazar et al., 2023), as well as detecting models that have already been trained on evaluation data (Carlini et al., 2018; Jagielski, 2023; Marone & Van Durme, 2024). Jacovi et al. (2023) on the other hand, suggest three strategies for reducing the risks of contamination. Data decontamination is an important step in ensuring the integrity of model evaluation, and is included in the discussion section because this critical step is relevant, and should be required, for all stages of model training.
@@
Test set contamination methods often follow a similar form to deduplication methods. When decontaminating MassiveTextRae et al. (2021) compute the 13-gram Jaccard similarity between train and test documents (as originally proposed by Lee et al. (2022a)) and remove train documents that have a Jaccard similarity over 0.8, the same exact method that they use to detect duplicated documents. Similarly, Brown et al. (2020) and Gao et al. (2020) use 13-gram overlap filtering to detect contamination in their respective training sets. To make decontamination easier, some tools have been developed. For example, CarperAI create a GitHub repository with code for decontaminating evaluation data from a training dataset using MinHashLSH. [^20] CarperAI decontamination GitHub repository Additionally, Marone & Van Durme (2024) propose Data Portraits, a method of investigating the data that a model was trained on. Data portraits are an artifact that records training data, allows downstream inspection, and enables answering questions about test set leakage and model plagiarism through a novel membership testing tool, a strided bloom filter. In addition to the new method, they release a demo of their tool [^21] https://dataportraits.org/ which uses 50-character grams to detect overlap between training data and an input sequence.
@@
Proving that black-box models have trained on a data point can be very challenging, but also important to ensure that their results can be trusted. Because their data is not disclosed, it is impossible to know with certainty what their training data was, but some methods have been proposed as efforts to address this issue. Shi et al. (2024) propose min-kprob, using the intuition that unseen examples are more likely to contain outlier words with low probabilities according to the model, whereas unseen examples are less likely to exhibit such low probability words. While this is not explicitly a method of decontamination, it is a useful tool for determining how contaminated a pretrained models data was. This tool can be used for benchmark example contamination detection, privacy auditing for machine unlearning, and copyrighted text detection. Oren et al. (2024) present another method for detecting contaminated data based on the observation that models have a tendency to memorize the order of examples as they exist in the dataset, and provide provable guarantees of test set contamination through a framework of statistical tests.
@@
Decontamination in code data is particularly important, as Li et al. (2022a) suggest that competitors in coding competitions tend to publish their solutions online after a competition. One solution to this problem is to use a temporal split when dividing data into training and evaluation. To ensure that test data does not exist in the training dataset, the training dataset can be constructed only from files that originate prior to the test competition. This ensures that the training data includes only information that would be available for human participants of the competition. In a more traditional method, StarCoder (Li et al., 2023b) decontaminates The Stack by removing files that contained either exact matching docstrings or solutions from the HumanEval (Chen et al., 2021) or MBPP (Austin et al., 2021) datasets, docstrings from APPS (Hendrycks et al., 2021), questions from GSM8K (Cobbe et al., 2021), or prompts from DS1000 (Lai et al., 2023).
@@ 10.4 Tools for data selection
For a more comprehensive and up-to-date list of available tools, see the Foundation Model Development Cheatsheet (Longpre et al., 2024).
@@ 11.1 Accelerating research on data selection
Research on data can be a very slow and tedious process, potentially requiring significant investments of time, compute, and human effort. Here we describe four directions of future research that can reduce the effort required, and the barriers to entry, for research in data selection.
Scaling down.
One direction that can enable more researchers to participate, as well as allowing for faster iteration is scaling down the current dataset sizes and models. For example, Kaddour (2023) demonstrate that it is possible to train a model on 745x less data, but lose only 1.9% performance on GLUE (Wang et al., 2018a) and 2.5% performance on super-natural instructions (Wang et al., 2022b). Additionally, Polo et al. (2024) show that a wide variety of evaluation benchmarks can be distilled into only 100 data points, while maintaining an accuracy within  2% of the true accuracy. One area that is still uncertain is whether the results from small models scale up to larger models, or if there is any way to predict the results on large models, based on experiments from small models. This is an active area of research.
@@
Preliminary works have been done in this direction, including some works on data attribution and valuation (Ghorbani & Zou, 2019; Ilyas et al., 2022), and data measurements (Mitchell et al., 2022). Works on data valuation and attribution (discussed in §9) attempt to better understand how each data point affects a models predictions, but may be used as inspiration for methods that directly evaluate the data itself. Additionally, there are a number of metrics that directly measure the intrinsic characteristics of data including distance, density, diversity, tendencies (mean, median, mode), and association (correlation, mutual information), but these measures have yet to be connected with a model’s performance on downstream data (Mitchell et al., 2022).
@@ 11.3 Shifting compute time from model training to data processing
One example of how this may be accomplished is through the use of methods which have previously only been used in low-data settings (e.g. alignment) and applying them to settings with greater amounts of data (e.g. pretraining). For example, methods inspired by model-based filtering, or reward model-reweighting techniques can be used to improve the “quality” of pretraining data. Additionally, techniques for generating synthetic data can be used to create more diverse training data Bradley et al. (2023) and methods previously used only for red-teaming (Samvelyan et al., 2024) can be used to generate training data for more robust models. As compute gets cheaper, it becomes more feasible to spend the cost of directly applying the existing methods. However, another avenue worth exploring is using cheap approximations for utility functions, such as computing similarity using bag-of-n-grams and hash functions rather than from model representations.
@@ Acknowledgments
This material is based on work that is partially funded by an unrestricted gift from Google. This work was supported by the National Science Foundation award #2048122. The views expressed are those of the authors and do not reflect the official policy or position of the US government. The work of BH and SC is partially supported by the UC Santa Barbara IEE IGSB SW Impact Grant. We want to thank Luca Soldaini for sharing their figure that was adapted into Figure 4. We are also thankful to Lucy Li for sharing her summary of different filtering mechanisms used for language models of recent years. Some of the figures include icons from the Noun Project: https://thenounproject.com/.